OpenStudy (anonymous):

f(x)=x^2-2x-8 find the vertex x-int and y-int

myininaya (myininaya):

f(x)=(x^2-2x+1)-8-1=(x-1)^2-9 Vertex is (1,-9) find x intercept let y=0 solve for x find y intercept let x=0 solve for y

myininaya (myininaya):

0=(x-1)^2-9 solve this for x y=(0-1)^2-9 solve this for y